when i drive to work it takes me 35 minutes. when i take the train only 15 minutes. what is the distance to work if the train is traveling an average of 80mph faster than i drive?
Answer by richwmiller(17219)

(35/60)x=1/4*(x+80)
x=60
x+80=140
35 miles